OpenStudy (anonymous):

If you have an equation like 3 = (1 + x/12)^12 How do you solve for x ? Like how do you get rid of that pesky exponent?

OpenStudy (anonymous):

First, take the 12th root of both sides (raise both sides to the power of 1/12). Why? A property of exponents is that if you have a term with an exponent in it to another exponent, you multiply them together. Since words are confusing, here's an example: (x^12)^(1/12) = x^(12 * 1/12) = x (As you can see, I conveniently used 12 since your problem has an exponent of 12. Going back to the problem after everything is raised to the power of 1/12, we have: 3^(1/12) = 1 + x/12 Doing basic algebra to arrive at the solution now: (3^(1/12)) -1 = x/12 12((3^(1/12)) - 1 ) = x x is equal to around 1.15.

OpenStudy (anonymous):

thank you.. thats what I was missing. I was thinking there must be some way to invert the ^12 on both sides, but wasn't clicking thought it might require a log of some kind.

OpenStudy (anonymous):

logs also work.

OpenStudy (agent0smith):

logs would make things more complicated.

OpenStudy (agent0smith):

If\[\Large x^n = a\]then take the nth root of both sides\[\Large x = \sqrt[n]{a}\]